OpenStudy (anonymous):

- 5/17p= - 8 how do i solve this step by step?

OpenStudy (anonymous):

Multiply the reciprocal of the fraction with the variable to both sides. It should look like this: (-17/5)-5/17p = -8/1(-17/5) Cancel out (-17/5)-5/17. This leaves you with p = -8/1(-17/5). Multiply -8/1(-17/5) and you will get 136/5. Turn it into a mixed fraction. p = 27 1/5 Hopefully this helps, I'm not the best explainer.

OpenStudy (anonymous):

Is it -(5/17)p OR -5/(17p)?